OpenStudy (anonymous):

solve 2/x-1 + 4/x = 8/x(x-1)

OpenStudy (anonymous):

x(2/x-1+4/x=x(8/x)(x-1) 2-x+4=8x-8 14=7x x=2

OpenStudy (anonymous):

can you explain please

OpenStudy (anonymous):

\[\frac{ 2 }{ x-1 }+\frac{ 4 }{ x }=\frac{ 8 }{ x(x+1) }\] like this?

OpenStudy (anonymous):

yes like that

OpenStudy (anonymous):

\[(\frac{ 2 }{ x-1 }+\frac{ 4 }{ x }=\frac{ 8 }{ (x)(x-1) })x(x-1)\] like @spiel7 said but it come out a little different then he said

OpenStudy (anonymous):

what do you get when you do this?

OpenStudy (anonymous):

i saw the equation incorrectly. big teach has the ball. find the lcd and you are home

OpenStudy (anonymous):

sorry i dont really get it

OpenStudy (anonymous):

common lcd is x(x+1)(x-1)

OpenStudy (anonymous):

once you do these it will get rid of all the fractions and you will be able to solve the normal way all we are doing is multippling by all the denomenators

OpenStudy (anonymous):

there isn't an x+1 denominator, it's x-1

OpenStudy (anonymous):

sorry i missed typed the first ime then. So do not multiply by the x+1 just x and x-1 this will make it a lot easier :)

OpenStudy (anonymous):

yyou should get \[2x+4(x-1)=8\] do you see how?

OpenStudy (anonymous):

yea i get it

OpenStudy (anonymous):

cool sorry about the confusion!! :) So get how to finish the equation

OpenStudy (anonymous):

so 2x+ 4x-4=8 move over the 8 then factor

OpenStudy (anonymous):

there are no x^2 in here so no factor easy solve

OpenStudy (anonymous):

6x-4=8

OpenStudy (anonymous):

oh yea yea i dont factor, ok yea sorry got it

OpenStudy (anonymous):

np

OpenStudy (anonymous):

thanks